OpenStudy (anonymous):

A tuition scale for a state university shows an increase of $1,500 each year. If the initial cost was $23,000. What will the cost of tuition be in twenty five years? In complete sentences, explain how an arithmetic sequence was applied?

OpenStudy (anonymous):

Total=23,000+1500x x is number of years so Total=23,000+1500(25) Total=60,500

OpenStudy (anonymous):

Arithmetic sequence was applied because between every year there is a $1500 difference